Algebraically sum the two electric field waves

E1=E01cos([omega]t + [alpha]1) and E2=E02([omega]t +[alpha]2)

under the condtions E02=E01 and aplha1=0. Your result should be a single harmonic (sine or cosine) function with appropriate amplitude and phase. Express the amplitude and phase in terms of parameters of the two combined waves.
Answer #1

Electric fields,

E1=E01\cos(\omega t+\alpha _{1})

And,

E2=E02\cos(\omega t+\alpha _{2})

The angular frequency of the two electric field are same.

The amplitude of the two electric fields are,

E_{sum}=\sqrt{(E01)^{2}+(E02)^{2}}

It is given that,

E02=E01

So, the sum of two electric fields amplitude are,

E_{sum}=\sqrt{(E01)^{2}+(E01)^{2}}

E_{sum}=\sqrt{2E01^{2}}

{\color{Blue} E_{sum}=\sqrt{2}E01}

The value of phase is,

\alpha=\alpha _{1}+\alpha _{2}

it is given that,

\alpha _{1}=0

therefore, the value of phase angle is,

\alpha =0+\alpha _{2}

{\color{Blue} \alpha =\alpha _{2}}

Express the amplitude and phase in terms of parmeters of the two combined waves is,

{\color{Blue} E_{sum}=\sqrt{2}E01(\cos(\omega t+\alpha 2))}